OpenStudy (anonymous):

Which of the following is the simplified form of y^2+y-12/y^2-9 ? a. , with the restriction y ≠ -3 b. , with the restriction y ≠ 3 c. , with the restriction y ≠ ±3 d. , with the restriction y ≠ ±3

OpenStudy (anonymous):

Please just give me the letter, Ive been working on this for hours :(

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what's the difference between C & D

OpenStudy (anonymous):

\[\frac{ y ^{2}+y-12 }{ y ^{2}-9 }\]

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

y+4/y+3 , with the restriction y ≠ -3 y-4/y-3 , with the restriction y ≠ 3 y+4/y+3 with the restriction y ≠ ±3 y-4/y-3 with the restriction y ≠ ±3

OpenStudy (anonymous):

those are the options

OpenStudy (anonymous):

\[\frac{ (y-3)\times(y-4) }{ (y-3)(y+3) }\]

OpenStudy (anonymous):

excuse me

OpenStudy (anonymous):

\[\frac{ (y-3)\times(y+4)}{ (y-3)\times(y+3) }\]

OpenStudy (anonymous):

can you just give me the answer you have no idea how frustrated i am, ive been working on this for hours

OpenStudy (anonymous):

it's a)

OpenStudy (anonymous):

it's Ok